solve: rex can paint a motel room in 3h and carl can do it in 6h. they work together for 1h. how long will it take rex to finish the job alone?
please help me. i've tried to get this and i can't understand.
thank you so much!

Answer by math_helper(2461)


solve: rex can paint a motel room in 3h and carl can do it in 6h. they work together for 1h. how long will it take rex to finish the job alone?
please help me. i've tried to get this and i can't understand.
thank you so much!
=================================================================
The key is to determine how much work they each do per unit of time.
Rex: 1/3 of a room per hour
Carl: 1/6 of a room per hour
Together: (1/3) + (1/6) = (2/6) + (1/6) = 3/6 = 1/2 of a room per hour
They work together for 1hr: that's 1/2 of the room.
So how long will it take Rex to finish the remaining 1/2 of the room?
(Amount of work)/(rate of work) = time to complete that work
(1/2 room) / (1/3 room per hour) = ? hours
——